# Temperature is All You Need for Generalization in Langevin Dynamics and other Markov Processes

### v1

@@ A.2 General Lemmas: Data processing inequality and generalized second laws of thermodynamics
The first direction follows directly from Claim 2.3, so we only need to prove the converse. Assume that either $\mathrm{KL}\left(p\,\middle\|\,q\right)+\mathrm{KL}\left(q\,\middle\|\,p\right)\leq\beta$, or $D_{\infty}\left(p\,\|\,q\right)+D_{\infty}\left(q\,\|\,p\right)\leq\beta$. In these cases, both $\mathop{}\!\mathrm{d}p/\mathop{}\!\mathrm{d}q$ and $\mathop{}\!\mathrm{d}q/\mathop{}\!\mathrm{d}p$ exist, and for any measurable event $B$, $p\left(B\right)=0\iff q\left(B\right)=0$, or equivalently, $p\left(B\right)>0\iff q\left(B\right)>0$. Therefore, $\mathrm{supp}\left(p\right)=\mathrm{supp}\left(q\right)$, and $\mathop{}\!\mathrm{d}p/\mathop{}\!\mathrm{d}q>0$ on $\mathrm{supp}\left(p\right)$. Denote $\Psi=-\ln\mathop{}\!\mathrm{d}p/\mathop{}\!\mathrm{d}q$, then $p$ is Gibbs w.r.t. $q$ with potential $\Psi$. The same derivation as in the proof of Claim 2.3 results in the bounds $\mathbb{E}_{q}\Psi-\mathbb{E}_{p}\Psi\leq\beta$ and $\operatorname{ess\,sup}_{q}\Psi-\operatorname{ess\,inf}_{p}\Psi\leq\beta$. In particular, if the latter holds then $\Psi$ can be shifted such that essentially $0\leq\Psi\leq\beta$. ∎
@@ Appendix C Tightness and Necessity of the Divergence Conditions
If we are only interested in ensuring generalization at time $t\rightarrow\infty$, and when we converge to the stationary distribution $p_{\infty}$, then it is enough to bound the divergence $D\left(p_{\infty}\,\|\,\nu\right)$. If we are interested in bounding $D\left(p_{t}\,\|\,\nu\right)$ (and consequently, the generalization gap) at all times $t$, then we need also to limit $p_{0}$’s dependence on $S$, since $p_{0}$ (as well as $p_{t}$ for small $t$) can be completely different from a stationary $p_{\infty}$, and just bounding $D\left(p_{\infty}\,\|\,\nu\right)$ does not say anything about it. Bounding $D\left(p_{0}\,\|\,\mu\right)$, for some data-independent distribution $\mu$, ensures generalization at $p_{0}$. This leaves the following questions regarding the proof of Theorem 2.7:
  • Why do we need to bound the divergences $D\left(p_{\infty}\,\|\,\nu\right)$ and $D\left(p_{0}\,\|\,\nu\right)$ from the same distribution $\nu$? That is, require $\mu=\nu$?
  • Why do we need to also bound the reverse divergence $D\left(\nu\,\|\,p_{\infty}\right)$? I.e., why do we need to require $p_{\infty}$ is Gibbs w.r.t. $\nu$ with a bounded potential, instead of just controlling the divergence $D\left(p_{\infty}\,\|\,\nu\right)$?
As we now show, without either one of these we cannot ensure generalization at intermediate times $t\geq 0$.
@@
Appendix D Generalized Version of Corollary 3.1
We start by characterizing the stationary distributions of SDERs in a box with different noise scales $\sigma^{2}$. The stationary distributions for Gaussian initialization can be found similarly. Then, we extend Corollary 3.1 to scenarios where $p_{0}\neq\nu$, as an immediate consequence of Theorem 2.7.
@@ D.1 Stationary distributions of CLD
with normal reflection in a box domain (for a full definition see (41)-(43) in Section H.2), where $L\geq 0$ is some $\mathcal{C}^{2}$ loss function, $\beta>0$ is an inverse temperature parameter, and $\sigma^{2}$ is a diffusion coefficient. First, we present a well known characterization of the stationary distribution of (21).
@@
If $L,\sigma^{2}\in\mathcal{C}^{2}$, $\sigma^{2}\left(\cdot\right)>0$ is uniformly bounded away from 00 in $\overline{\Omega}$,
[…]
and the integrals exist, then
[…]
@@
For completeness, the proof as well as additional definitions are presented in Section H.1. We can now calculate explicit stationary distributions for some choices of $\sigma^{2}$.
@@ D.2 Generalization bounds
Assume that training follows a CLD in a bounded domain as described in (21) with uniform initialization $p_{0}=\mathrm{Uniform}\left(\Theta_{0}\right)$, where $\Theta_{0}\subseteq\Theta$. For simplicity we take $\sigma^{2}\equiv 1$. In that case Theorem 2.7 implies the following.
@@
    $\mathbb{E}_{{\boldsymbol{\theta}}_{t}\sim p_{t}}\left[E_{D}({\boldsymbol{\theta}}_{t})-E_{S}({\boldsymbol{\theta}}_{t})\right]\leq\sqrt{\frac{\beta\mathbb{E}_{p_{0}}L_{S}({\boldsymbol{\theta}})+\ln{\left\lvert{\Theta}\right\rvert}/{\left\lvert{\Theta_{0}}\right\rvert}+\ln\left(1/\delta\right)}{2N}}\,.$ (28)
@@
    $E_{D}({\boldsymbol{\theta}}_{t})-E_{S}({\boldsymbol{\theta}}_{t})\leq\sqrt{\frac{\beta\operatorname{ess\,sup}_{p_{0}}L_{S}({\boldsymbol{\theta}})+\ln{\left\lvert{\Theta}\right\rvert}/{\left\lvert{\Theta_{0}}\right\rvert}+\ln\left(1/\delta\right)}{2N}}\,.$ (29)
@@
This is a direct corollary of Theorem 2.7 with $\mathrm{KL}\left(p_{0}\,\middle\|\,\nu\right)=\ln{\left\lvert{\Theta}\right\rvert}/{\left\lvert{\Theta_{0}}\right\rvert}$. ∎
@@
Let ${\boldsymbol{\lambda}}\in\mathbb{R}^{d}_{>0}$ and consider the unconstrained SDE
@@
Notice that $-\beta^{-1}\operatorname{diag}\left({\boldsymbol{\lambda}}\right){\boldsymbol{\theta}}_{t}dt$ corresponds to an additive regularization of the form $\frac{1}{2\beta}{\boldsymbol{\theta}}_{t}^{\top}\operatorname{diag}\left({\boldsymbol{\lambda}}\right){\boldsymbol{\theta}}_{t}$, so each parameter can have a different regularization coefficient. We shall denote by $\phi_{{\boldsymbol{\lambda}}}$ a multivariate Gaussian distribution with mean $\mathbf{0}$ and covariance matrix $\operatorname{diag}\left({\boldsymbol{\lambda}}^{-1}\right)$, where ${\boldsymbol{\lambda}}^{-1}=\left(\lambda_{1}^{-1},\dots,\lambda_{d}^{-1}\right)$. For simplicity, we present the results with $\sigma^{2}\equiv 1$.
@@
Let ${\boldsymbol{\lambda}}_{0},{\boldsymbol{\lambda}}_{1}>0$, and let ${\boldsymbol{\theta}}_{t}$ evolve according to (30) with $\sigma^{2}\equiv 1$, ${\boldsymbol{\lambda}}={\boldsymbol{\lambda}}_{1}$. Then for any time $t\geq 0$, and $\delta\in\left(0,1\right)$,
@@
    $\mathbb{E}_{{\boldsymbol{\theta}}_{t}\sim p_{t}}\left[E_{D}({\boldsymbol{\theta}}_{t})-E_{S}({\boldsymbol{\theta}}_{t})\right]\leq\sqrt{\frac{\beta\mathbb{E}_{p_{0}}L_{S}({\boldsymbol{\theta}})+\mathrm{KL}\left(\phi_{{\boldsymbol{\lambda}}_{0}}\,\middle\|\,\phi_{{\boldsymbol{\lambda}}_{1}}\right)+\ln\left(1/\delta\right)}{2N}}\,.$ (31)
@@
    $E_{D}({\boldsymbol{\theta}}_{t})-E_{S}({\boldsymbol{\theta}}_{t})\leq\sqrt{\frac{\beta\operatorname{ess\,sup}_{p_{0}}L_{S}({\boldsymbol{\theta}})+\mathrm{KL}\left(\phi_{{\boldsymbol{\lambda}}_{0}}\,\middle\|\,\phi_{{\boldsymbol{\lambda}}_{1}}\right)+\ln\left(1/\delta\right)}{2N}}\,,$ (32)
@@
This is a direct corollary of Theorem 2.7 with the explicit expression for the KL divergence between two Gaussians. ∎
@@
While the bound in Lemma D.7 depends on the dimension of the parameters $d$, this can be mitigated in practice. For example, by matching the regularization coefficient and initialization variance, the KL-divergence term vanishes and we lose the dependence on dimension. Furthermore, we can control each parameter separately by using parameter specific initialization variances and regularization coefficients. Then, the KL-divergence can have different dependencies, if any, on the dimension $d$.
@@ Appendix E Linear Regression with CLD
Theorems 2.7 and 3.1 only bound the gap between the population and training errors, yet this does not necessarily bound the population error itself. One way to do this is by separately bounding the training error and showing that in the regime in which the generalization gap is small, the training error can be small as well. In Appendix F we show empirically that deep NNs can reach low training error when trained with SGLD in the regime in which Corollary 3.1 is not vacuous. Here, we look at the particular case of the asymptotic behavior of ridge regression with CLD training with Gaussian i.i.d. data, for which we can analytically study the training and population losses.
@@
In the rest of this section, our final aim is to calculate the expected training and population losses in the setup described above, in the case when the data is sampled i.i.d. from standard Gaussian distribution, $\sigma$ is a fixed constant, $\lambda\propto d$ (to match standard initialization [^5] Since this is a linear model $d=\mathrm{layer\,width}$, and as we assume the regularization matches the standard initialization. ), $N,\beta$ and $d$ are large, but $\beta\ll N$, so our generalization bound is small (since $\mathbb{E}_{p_{0}}L$ is a fixed constant in this case). We will find (in Remark E.2 and Remark E.4) that if also $d\ll\beta$ then the training and expected population loss are not significantly degraded. This is not a major constraint, since we need $d\ll N$ to get good population loss anyway, even without noise (i.e. $\beta\rightarrow\infty$). This shows that in this regime $d\ll\beta\ll N$, the randomness required by our generalization bound (the KL bounds in Corollary 3.1) does not significantly harm the training loss or the expected population loss.
@@
We intuitively derive the asymptotic behavior of Claim E.1. Let $\lambda$ be constant, and let $\beta$ grow (so $\alpha$ shrinks). We can decompose $\left(\mathbf{A}+\alpha\mathbf{I}\right)^{-1}$ as
@@
then $\left\|R_{2}\left(\alpha\right)\right\|_{2}\leq\frac{\alpha^{2}}{\lambda_{\min}\left(\mathbf{A}\right)^{3}}$, where $\lambda_{\min}\left(\mathbf{A}\right)$ is the minimal eigenvalue of $\mathbf{A}$. As the elements of $\mathbf{X}$ are i.i.d. with mean 00 and variance $1$, the limiting distribution of the spectrum of $\mathbf{A}$ as $N,d\to\infty$ with $d/N\to\gamma\in\left(0,1\right)$ is the Marchenko–Pastur distribution, which is supported on $\left[\left(1-\sqrt{\gamma}\right)^{2},\left(1+\sqrt{\gamma}\right)^{2}\right]$. In particular, as $N,d\to\infty$, $\lambda_{\min}\left(\mathbf{A}\right)\geq\left(1-\sqrt{d/N}\right)^{2}$, so for $\varepsilon>0$,
@@
Therefore, the added noise does not significantly hurt the training loss when $\frac{1}{\beta}\lessapprox\sigma^{2}\left(\frac{1}{d}-\frac{1}{N}\right)$, or equivalently, $\beta\gtrapprox\frac{Nd}{(N-d)\sigma^{2}}$. In particular, this holds when $d\ll\beta\ll N$, which is a regime where our generalization bound Corollary 3.1 also becomes small (since $\beta\ll N$). This shows that the randomness required by Corollary 3.1 can allow for successful optimization of the training loss.
@@
As we have done for the training loss in Remark E.2, we can estimate the expected population loss in some asymptotic regimes. Let $\lambda$ be constant, and let $\beta$ grow (so $\alpha$ shrinks). As in Remark E.2, we use the approximation $\left(\mathbf{A}+\alpha\mathbf{I}\right)^{-1}=\mathbf{A}^{-1}-\alpha\mathbf{A}^{-2}+O\left(\alpha^{2}\mathbf{I}\right)$, which also implies $\left(\mathbf{A}+\alpha\mathbf{I}\right)^{-2}=\mathbf{A}^{-2}-2\alpha\mathbf{A}^{-3}+O\left(\alpha^{2}\mathbf{I}\right)$, and treat the remainders as $O\left(\alpha^{2}\right)$ even when taking the expectation w.r.t. $\mathbf{X}$. Then,
@@
This result is similar to the one in Remark E.2 — for the expected population loss not to be significantly hurt by the added noise, it must hold that $\beta\gtrapprox\frac{Nd}{\left(N-1\right)\sigma^{2}}$. In particular, this holds when $d\ll\beta\ll N$, which is a regime where our generalization bound Corollary 3.1 also becomes small (since $\beta\ll N$). This shows that the randomness required by Corollary 3.1 does not harm the expected population loss.

### v2

@@ B.3 Single-Sample PAC-Bayes Bounds
The first direction follows directly from Claim˜2.3, so we only need to prove the converse. Assume that either $\mathrm{KL}\left(p\,\middle\|\,q\right)+\mathrm{KL}\left(q\,\middle\|\,p\right)\leq\beta$, or $D_{\infty}\left(p\,\|\,q\right)+D_{\infty}\left(q\,\|\,p\right)\leq\beta$. In these cases, both $\mathop{}\!\mathrm{d}p/\mathop{}\!\mathrm{d}q$ and $\mathop{}\!\mathrm{d}q/\mathop{}\!\mathrm{d}p$ exist, and for any measurable event $B$, $p\left(B\right)=0\iff q\left(B\right)=0$, or equivalently, $p\left(B\right)>0\iff q\left(B\right)>0$. Therefore, $\mathrm{supp}\left(p\right)=\mathrm{supp}\left(q\right)$, and $\mathop{}\!\mathrm{d}p/\mathop{}\!\mathrm{d}q>0$ on $\mathrm{supp}\left(p\right)$. Denote $\Psi=-\ln\mathop{}\!\mathrm{d}p/\mathop{}\!\mathrm{d}q$, then $p$ is Gibbs w.r.t. $q$ with potential $\Psi$. The same derivation as in the proof of Claim˜2.3 results in the bounds $\mathbb{E}_{q}\Psi-\mathbb{E}_{p}\Psi\leq\beta$ and $\operatorname{ess\,sup}_{q}\Psi-\operatorname{ess\,inf}_{p}\Psi\leq\beta$. In particular, if the latter holds then $\Psi$ can be shifted such that essentially $0\leq\Psi\leq\beta$. ∎
@@ Appendix C Tightness and Necessity of the Divergence Conditions
If we are only interested in ensuring generalization at time $t\rightarrow\infty$, and when we converge to the stationary distribution $p_{\infty}$, then it is enough to bound the divergence $D\left(p_{\infty}\,\|\,\nu\right)$. If we are interested in bounding $D\left(p_{t}\,\|\,\nu\right)$ (and consequently, the generalization gap) at all times $t$, then we need also to limit $p_{0}$’s dependence on $S$, since $p_{0}$ (as well as $p_{t}$ for small $t$) can be completely different from a stationary $p_{\infty}$, and just bounding $D\left(p_{\infty}\,\|\,\nu\right)$ does not say anything about it. Bounding $D\left(p_{0}\,\|\,\mu\right)$, for some data-independent distribution $\mu$, ensures generalization at $p_{0}$. This leaves the following questions regarding the proof of Theorem˜2.7:
  • Why do we need to bound the divergences $D\left(p_{\infty}\,\|\,\nu\right)$ and $D\left(p_{0}\,\|\,\nu\right)$ from the same distribution $\nu$? That is, we do we need to require $\mu=\nu$? Bounding the divergences of $p_{0}$ and $p_{\infty}$ to two different divergences $\mu\neq\nu$ is sufficient to get generalization at the beginning (i.e. initialization) and end (i.e. after mixing)–is it sufficient for generalization in the middle (i.e. at any $t$)?
  • Why do we need to also bound the reverse divergence $D\left(\nu\,\|\,p_{\infty}\right)$? I.e., why do we need to require $p_{\infty}$ is Gibbs w.r.t. $\nu$ with a bounded potential, instead of just controlling the divergence $D\left(p_{\infty}\,\|\,\nu\right)$, which is a weaker requirement and sufficient for generalization after mixing?
As we now show, both are necesairy, and without requiring both, i.e. if we drop either one of these, we cannot ensure generalization at intermediate times $t\geq 0$.
@@
Appendix D Generalized Version of Corollary˜3.1
We start by characterizing the stationary distributions of SDERs in a box with different noise scales $\sigma^{2}$. The stationary distributions for Gaussian initialization can be found similarly. Then, we extend Corollary˜3.1 to scenarios where $p_{0}\neq\nu$, as an immediate consequence of Theorem˜2.7.
@@ D.1 Stationary distributions of CLD
with normal reflection in a box domain (for a full definition see (45)-(47) in Section˜H.2), where $L\geq 0$ is some $\mathcal{C}^{2}$ loss function, $\beta>0$ is an inverse temperature parameter, and $\sigma^{2}$ is a diffusion coefficient. First, we present a well known characterization of the stationary distribution of (25).
@@
If $L,\sigma^{2}\in\mathcal{C}^{2}$, $\sigma^{2}\left(\cdot\right)>0$ is uniformly bounded away from 0 in $\overline{\Omega}$,
$Z=\intop_{\overline{\Omega}}\frac{1}{\sigma^{2}\left(\mathbf{x}\right)}\exp\left(-\beta\intop\frac{\nabla L\left(\mathbf{x}\right)}{\sigma^{2}\left(\mathbf{x}\right)}d\mathbf{x}\right)<\infty\,$,
the integrals exist, and the field $\nabla L/\sigma^{2}$ is conservative (curl-free), then
$p_{\infty}\left(\mathbf{x}\right)=\frac{1}{Z}\frac{1}{\sigma^{2}\left(\mathbf{x}\right)}\exp\left(-\beta\intop\frac{\nabla L\left(\mathbf{x}\right)}{\sigma^{2}\left(\mathbf{x}\right)}d\mathbf{x}\right)\,$ (26)
@@
For completeness, the proof is presented in Section˜H.2.1, following additional results and definitions in Appendix˜H. We can now calculate explicit stationary distributions for some choices of $\sigma^{2}$. Specifically, we focus on cases where $\sigma^{2}\left(\mathbf{x}\right)=g\left(L\left(\mathbf{x}\right)\right)$ for some scalar function $g$, as it guarantees the curl-free condition, and is convenient to integrate.
@@ D.2 Generalization bounds
Assume that training follows a CLD in a bounded domain as described in (25) with uniform initialization $p_{0}=\mathrm{Uniform}\left(\Theta_{0}\right)$, where $\Theta_{0}\subseteq\Theta$. For simplicity we take $\sigma^{2}\equiv 1$. In that case Theorem˜2.7 implies the following.
@@
    $\mathbb{E}_{{\boldsymbol{\theta}}_{t}\sim p_{t}}\left[E_{D}({\boldsymbol{\theta}}_{t})-E_{S}({\boldsymbol{\theta}}_{t})\mid S\right]\leq\sqrt{\frac{\beta\mathbb{E}_{p_{0}}\left[L_{S}({\boldsymbol{\theta}})\mid S\right]+\ln{\left\lvert{\Theta}\right\rvert}/{\left\lvert{\Theta_{0}}\right\rvert}+\ln\left(N/\delta\right)}{2N}}\,.$ (32)
@@
    $E_{D}({\boldsymbol{\theta}}_{t})-E_{S}({\boldsymbol{\theta}}_{t})\leq\sqrt{\frac{\beta\operatorname{ess\,sup}_{p_{0}}L_{S}({\boldsymbol{\theta}})+\ln{\left\lvert{\Theta}\right\rvert}/{\left\lvert{\Theta_{0}}\right\rvert}+\ln\left(N/\delta\right)}{2N}}\,.$ (33)
@@
This is a direct corollary of Theorem˜2.7 with $\mathrm{KL}\left(p_{0}\,\middle\|\,\nu\right)=\ln{\left\lvert{\Theta}\right\rvert}/{\left\lvert{\Theta_{0}}\right\rvert}$. ∎
@@
Let ${\boldsymbol{\lambda}}\in\mathbb{R}^{d}_{>0}$ be regularization terms, and consider the unconstrained SDE
@@
Notice that $\beta^{-1}\operatorname{diag}\left({\boldsymbol{\lambda}}\right){\boldsymbol{\theta}}_{t}dt$ corresponds to an additive regularization of the form $\frac{1}{2\beta}{\boldsymbol{\theta}}_{t}^{\top}\operatorname{diag}\left({\boldsymbol{\lambda}}\right){\boldsymbol{\theta}}_{t}$, so each parameter can have a different regularization coefficient. We shall denote by $\phi_{{\boldsymbol{\lambda}}}$ a multivariate Gaussian distribution with mean $\mathbf{0}$ and covariance matrix $\operatorname{diag}\left({\boldsymbol{\lambda}}^{-1}\right)$, where ${\boldsymbol{\lambda}}^{-1}=\left(\lambda_{1}^{-1},\dots,\lambda_{d}^{-1}\right)$. For simplicity, we present the results with $\sigma^{2}\equiv 1$.
@@
Let ${\boldsymbol{\lambda}}_{0},{\boldsymbol{\lambda}}_{1}>0$, and let ${\boldsymbol{\theta}}_{t}$ evolve according to (34) with $\sigma^{2}\equiv 1$ and ${\boldsymbol{\lambda}}={\boldsymbol{\lambda}}_{1}$, and start from a Gaussian initialization $p_{0}=\phi_{{\boldsymbol{\lambda}}_{0}}$. Then for any time $t\geq 0$, and $\delta\in\left(0,1\right)$,
@@
    $\mathbb{E}_{{\boldsymbol{\theta}}_{t}\sim p_{t}}\left[E_{D}({\boldsymbol{\theta}}_{t})-E_{S}({\boldsymbol{\theta}}_{t})\mid S\right]\leq\sqrt{\frac{\beta\mathbb{E}_{p_{0}}\left[L_{S}({\boldsymbol{\theta}})\mid S\right]+\mathrm{KL}\left(\phi_{{\boldsymbol{\lambda}}_{0}}\,\middle\|\,\phi_{{\boldsymbol{\lambda}}_{1}}\right)+\ln\left(N/\delta\right)}{2N}}\,.$ (35)
@@
    $E_{D}({\boldsymbol{\theta}}_{t})-E_{S}({\boldsymbol{\theta}}_{t})\leq\sqrt{\frac{\beta\operatorname{ess\,sup}_{p_{0}}L_{S}({\boldsymbol{\theta}})+\mathrm{KL}\left(\phi_{{\boldsymbol{\lambda}}_{0}}\,\middle\|\,\phi_{{\boldsymbol{\lambda}}_{1}}\right)+\ln\left(N/\delta\right)}{2N}}\,$, (36)
@@
This is a direct corollary of Theorem˜2.7 with the explicit expression for the KL divergence between two Gaussians. ∎
@@
While the bound in Lemma˜D.7 depends on the dimension of the parameters $d$, this can be mitigated in practice. For example, by matching the regularization coefficient and initialization variance, the KL-divergence term vanishes and we lose the dependence on dimension. Furthermore, we can control each parameter separately by using parameter specific initialization variances and regularization coefficients. Then, the KL-divergence can have different dependencies, if any, on the dimension $d$.
@@ Appendix E Linear Regression with CLD
Theorems˜2.7 and 3.1 only bound the gap between the population and training errors, yet this does not necessarily bound the population error itself. One way to do this is by separately bounding the training error and showing that in the regime in which the generalization gap is small, the training error can be small as well. In Appendix˜F we show empirically that deep NNs can reach low training error when trained with SGLD in the regime in which Corollary˜3.1 is not vacuous. Here, we look at the particular case of the asymptotic behavior of ridge regression with CLD training with Gaussian i.i.d. data, for which we can analytically study the training and population losses.
@@
In the rest of this section, our final aim is to calculate the expected training and population losses in the setup described above, in the case when the data is sampled i.i.d. from standard Gaussian distribution, $\sigma$ is a fixed constant, $\lambda\propto d$ (to match standard initialization),  [^9] Since this is a linear model $d=\mathrm{layer\,width}$, and as we assume the regularization matches the standard initialization. This initialization is considered in many works as a Bayesian prior in various settings [35, 68]. $N,\beta$ and $d$ are large, but $\beta\ll N$, so our generalization bound is small (since $\mathbb{E}_{p_{0}}L$ is a fixed constant in this case). We will find (in Remark˜E.2 and Remark˜E.4) that if also $d\ll\beta$ then the training and expected population loss are not significantly degraded. This is not a major constraint, since we need $d\ll N$ to get good population loss anyway, even without noise (i.e. $\beta\rightarrow\infty$). This shows that in this regime $d\ll\beta\ll N$, the randomness required by our generalization bound (the KL bounds in Corollary˜3.1) does not significantly harm the training loss or the expected population loss.
@@
We intuitively derive the asymptotic behavior of Claim˜E.1. Let $\lambda$ be constant, and let $\beta$ grow (so $\alpha$ shrinks). We can decompose $\left(\mathbf{A}+\alpha\mathbf{I}\right)^{-1}$ as
@@
then $\left\|R_{2}\left(\alpha\right)\right\|_{2}\leq\frac{\alpha^{2}}{\lambda_{\min}\left(\mathbf{A}\right)^{3}}$, where $\lambda_{\min}\left(\mathbf{A}\right)$ is the minimal eigenvalue of $\mathbf{A}$. As the elements of $\mathbf{X}$ are i.i.d. with mean 0 and variance $1$, the limiting distribution of the spectrum of $\mathbf{A}$ as $N,d\to\infty$ with $d/N\to\gamma\in\left(0,1\right)$ is the Marchenko–Pastur distribution, which is supported on $\left[\left(1-\sqrt{\gamma}\right)^{2},\left(1+\sqrt{\gamma}\right)^{2}\right]$. In particular, as $N,d\to\infty$, $\lambda_{\min}\left(\mathbf{A}\right)\geq\left(1-\sqrt{d/N}\right)^{2}$, so for $\varepsilon>0$,
@@
Therefore, the added noise does not significantly hurt the training loss when $\frac{1}{\beta}\lessapprox\sigma^{2}\left(\frac{1}{d}-\frac{1}{N}\right)$, or equivalently, $\beta\gtrapprox\frac{Nd}{(N-d)\sigma^{2}}$. In particular, this holds when $d\ll\beta\ll N$, which is a regime where our generalization bound Corollary˜3.1 also becomes small (since $\beta\ll N$). This shows that the randomness required by Corollary˜3.1 can allow for successful optimization of the training loss.
@@
As we have done for the training loss in Remark˜E.2, we can estimate the expected population loss in some asymptotic regimes. Let $\lambda$ be constant, and let $\beta$ grow (so $\alpha$ shrinks). As in Remark˜E.2, we use the approximation $\left(\mathbf{A}+\alpha\mathbf{I}\right)^{-1}=\mathbf{A}^{-1}-\alpha\mathbf{A}^{-2}+O\left(\alpha^{2}\mathbf{I}\right)$, which also implies $\left(\mathbf{A}+\alpha\mathbf{I}\right)^{-2}=\mathbf{A}^{-2}-2\alpha\mathbf{A}^{-3}+O\left(\alpha^{2}\mathbf{I}\right)$, and treat the remainders as $O\left(\alpha^{2}\right)$ even when taking the expectation w.r.t. $\mathbf{X}$. Then,
@@
This result is similar to the one in Remark˜E.2 — for the expected population loss not to be significantly hurt by the added noise, it must hold that $\beta\gtrapprox\frac{Nd}{\left(N-1\right)\sigma^{2}}$. In particular, this holds when $d\ll\beta\ll N$, which is a regime where our generalization bound Corollary˜3.1 also becomes small (since $\beta\ll N$). This shows that the randomness required by Corollary˜3.1 does not harm the expected population loss.